# SynthCoder: A Synthetical Strategy to Tune LLMs for Code Completion

### v2

@@ VI Threats to Validity
Limited datasets
The evaluation of SynthCoder is based on a limited number of datasets and benchmark methods, which may bring in certain validity risks. In particular, the test sets may not fully cover the diversity of real-world application scenarios, leading to potential gaps between benchmark results and practical performance. However, we have alreadly selected industry-recognized open-source benchmarks (ExecRepoEval and CrossCodeEval) and a custom test set constructed from third-party open-source datasets. Meanwhile, we evaluated the capabilities of SynthCoder and other baseline models in both single-file code completion, cross-file code completion based on dependency graph retrieval, and cross-file code completion based on BM25 retrieval so as to establish a comprehensive understanding of the performance of SynthCoderṪhe results strongly demonstrate the superiority of our method and model. Throughout the evaluation process, we utilized the official testing frameworks and parameters provided by the benchmarks or other well-known open-source models wherever possible to ensure the validity and generalizability of our tests.
Data leakage

### v3

@@ VI Threats to Validity
Data Leakage
Training Hyperparameter
Hyperparameters can significantly impact model performance. Due to computational constraints, we used the officially recommended settings for each benchmark model during comparisons rather than aligning hyperparameters across all methods. In this sense, we cannot entirely rule out the possibility that the performance gains we observed were solely attributable to the dataset we constructed. Nevertheless, we believe that the hyperparameter configurations used by other benchmark methods likely represented their respective optimal setups at the time. We encourage interested researchers to reproduce our work and explore potentially better hyperparameter configurations.
@@
Limited Benchmark Datasets
The evaluation of SynthCoder is based on a limited number of datasets and benchmark methods, which may bring in certain validity risks. In particular, the test sets may not fully cover the diversity of real-world application scenarios, leading to potential gaps between benchmark results and practical performance. However, we have alreadly selected industry-recognized open-source benchmarks. Meanwhile, we evaluated the capabilities of SynthCoder and other baseline models in both single-file code completion, cross-file code completion based on dependency graph retrieval, and cross-file code completion based on BM25 retrieval so as to establish a comprehensive understanding of the performance of SynthCoder . The results strongly demonstrate the superiority of our method and model. Throughout the evaluation process, we utilized the well-known testing frameworks and official parameters wherever possible to ensure the validity and generalizability of our tests.
@@ VII Conclusion
Building on our findings, we identify several promising directions for future research. Future work will advance beyond FIM towards more sophisticated edit prediction by learning from real-world code changes [50]. The practical deployment of such models hinges on a synergistic approach: creating compact, specialized models via techniques like knowledge distillation [51], and coupling them with system-level optimizations. This includes leveraging speculative decoding for low latency [52], paving the way for faster, smarter, and more dependable code assistants.